### v2

# A note on the Steinitz Lemma

Abstract.
We prove that the Euclidean Steinitz problem may be reduced to its restriction to “nearly unit” vectors at the cost of an additive $O\Big{(}\sqrt{\frac{d}{\log d}}\Big{)}$ error.
2020 Mathematics Subject Classification:
52A40, 46N10
Research of G. A. was partially supported by the ERC Advanced Grant "GeoScape" no. 882971, by Hungarian National Research (NKFIH) grants no. 147145, 147544, 150151, and KKP-133819, which has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the ADVANCED-24 funding scheme. Research of R. H. is supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2140004. Research was supported by project no. TKP2021-NVA-09, which has been implemented with the support provided by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021-NVA funding scheme.
1. Framework and results
More than a century ago, Steinitz [32] proved the following unexpected result.
Theorem 1 (Euclidean Steinitz lemma).
Given any finite family of vectors $V\subset\mathbb{R}^{d}$ of Euclidean norm at most 1 that sum to 0, one can order the elements of $V$ as $v_{1},\ldots,v_{n}$ so that for every $k=1,\ldots,n,$
(1) $\left|\sum_{i=1}^{k}v_{i}\right|\leq C$
where $C$ is a constant that depends only on the dimension $d$.
Here and later on, |.||.|| . | denotes the Euclidean norm on $\mathbb{R}^{d}$ whose unit ball is $B_{2}^{d}=\{x\in\mathbb{R}^{d}:|x|\leq 1\}$. More generally, $B_{p}^{d}$ will denote the unit ball of the $\ell_{p}$-norm on $\mathbb{R}^{d}$ for $1\leq p\leq\infty$. Also, for a finite set of vectors $V$, $\Sigma(V):=\Sigma_{v\in V}v$ will stand for the sum of the elements of $V$, and for $n\in\mathbb{N}_{+}$, we set $[n]:=\{1,\ldots,n\}$.
It is natural to ask for the smallest value of $C$ for which (1) is guaranteed to hold. We state the formal definition for general norms. Let $B\subseteq\mathbb{R}^{d}$ be a convex body, that is, a compact convex set with non-empty interior that satisfies $0\in\mathrm{int}\,B$, where $\mathrm{int}\,$ stands for interior. The Minkowski norm generated by $B$ is defined as
$\|x\|_{B}:=\inf\{r\geq 0:\ x\in rB\}$
(note that if $B$ is not symmetric about 00, then .$\|.\|_{B}$ is only positively homogeneous).
Definition 2 (Steinitz constant).
Let $B\subseteq\mathbb{R}^{d}$ be a convex body with $0\in\mathrm{int}\,B$. The Steinitz constant of $B$, denoted $S(B)$, is the smallest number $C$ for which any finite set of vectors $V\subset B$ with $\Sigma(V)=0$ has an ordering $V=\{v_{1},\ldots,v_{n}\}$ along which each partial sum has norm at most $C$. That is, for every $k\in[n]$,
(2) $\Big{\|}\sum_{i\in[k]}v_{i}\Big{\|}_{B}\leq C.$
The strongest general estimate on the Steinitz constants is as follows.
Theorem 3 (The Steinitz Lemma for general norms [32, 31, 17]).
For any convex body $B\subseteq\mathbb{R}^{d}$,
$S(B)\leq d.$
This theorem is tight for non-symmetric convex bodies, as is shown by taking $B$ to be the regular simplex centered at the origin and choosing $V$ to be the set of its vertices, whereas it is asymptotically sharp for symmetric norms, as is shown by the $\ell_{1}$ case [17]: $S(B_{1}^{d})\geq(d+1)/2$. An example in the Euclidean case [17] shows that $S(B_{2}^{d})\geq\sqrt{d+3}/2$ must hold.
The following long-standing conjecture of Bergström [9] would yield a much stronger estimate in the Euclidean norm:
Conjecture 4.
For all $d\geq 1$, $S(B_{2}^{d})=O(\sqrt{d})$.
We note that a matching bound is expected to hold for the maximum norm.
So far, Conjecture 4, known as the Steinitz problem, refuted all attempts. Our goal in this short note is to reduce the Euclidean Steinitz problem to its restriction to sets that consist of “nearly unit” vectors at the cost of an additive $O\Big{(}\sqrt{\frac{d}{\log d}}\Big{)}$ error. To that end, we introduce the following notion.
Definition 5 ($\varepsilon$-Steinitz constant).
For a fixed convex body $B\subseteq\mathbb{R}^{d}$ with $0\in\mathrm{int}\,B$, and for $0\leq\varepsilon\leq 1$, the $\varepsilon$-Steinitz constant of $B$, denoted $S_{\varepsilon}(B)$, is the smallest constant $C$ for which any finite set $V\subset\mathbb{R}^{d}$ consisting of vectors of .$\|.\|_{B}$-norm in $[1-\varepsilon,1]$ may be ordered as $V=\{v_{1},\ldots,v_{n}\}$ so that
(3) $\Big{\|}\sum_{i\in[k]}v_{i}-\frac{k}{n}\Sigma(V)\Big{\|}_{B}\leq C$
holds for every $k=1,\ldots,n$.
Note that condition $\Sigma(V)=0$ of Definition 2 is omitted and, consequently, (2) is replaced by inequality (3). This does not change the order of magnitude of $S_{\varepsilon}(B)$, as replacing each vector $v\in V$ by $v-\frac{1}{n}\Sigma(V)$ results in a family of vectors contained in $2B$ that sum up to 0. Furthermore, the estimate of Theorem 3 also holds in this more general setting (see [17]).
Clearly, $S_{1}(B_{2}^{d})=S(B_{2}^{d})$, and $S_{\varepsilon}(B_{2}^{d})\leq S(B_{2}^{d})$ for every $\varepsilon\in[0,1]$. Our main result establishes a reverse estimate.
Theorem 6.
For any $10<\varepsilon<$ 1 and all $d\geq 2$,
(4) $S(B_{2}^{d})\leq\frac{1}{\varepsilon}\Big{(}S_{\varepsilon}(B_{2}^{d})+200\sqrt{\frac{d}{\log d}}\Big{)}.$
In particular, an $o(d)$ estimate on the restricted problem would yield an $o(d)$ estimate for the original question: $S_{\varepsilon}(B_{2}^{d})=o(\varepsilon\cdot d)$ implies that $S(B_{2}^{d})=o(d)$. Moreover, it suffices to prove Conjecture 4 for vectors whose norm is in $[c,1]$ where $c<1$ is an arbitrary constant: $S(B_{2}^{d})=O(\sqrt{d})$ would follow from the estimate $S_{c}(B_{2}^{d})=O(\sqrt{d})$.
2. History of the Steinitz Lemma
Despite being more than a century old, the story of the Steinitz lemma is still far from complete. In the following, we provide an overview that lists the main, and often forgotten, steps in its development.
Every mathematician knows the Riemann series theorem, dating more than 150 years ago [28], which states that given a conditionally convergent series, one can rearrange it to converge to any real number. In other words, for any real series, the set of all sums of its rearrangements is either empty (meaning that the series is divergent), a single point (the series is absolutely convergent), or the entire real line. It is natural to ask for the extension of this result to series consisting of complex numbers and, even more generally, of vectors in $\mathbb{R}^{d}$. In his very first article in 1905, at the age of 19, Lévy [26] addressed this question and announced the following result.
Theorem 7 (Lévy-Steinitz theorem).
Given a series of vectors in $\mathbb{R}^{d}$, the set of all sums of its rearrangements is empty, or it forms an affine subspace of $\mathbb{R}^{d}$.
Recall that affine subspaces of $\mathbb{R}^{d}$ are of the form $L+x$ where $L\subset\mathbb{R}^{d}$ is a linear subspace and $x\in\mathbb{R}^{d}$.
Unfortunately, the proof of Lévy contained serious gaps for dimensions $d\geq 3$, as pointed out by Steinitz [32] in 1913. In turn, he gave the first complete proof of Theorem 7, known today as the Lévy-Steinitz theorem. Steinitz’s work is quite technical and wide-scoped: it was published in three parts  [32, 33, 34], with total length summing well over 100 pages. A key step of his proof is Theorem 1 (see [32, p.171]), which he stated with $C=2d$.
Independently of Lévy and Steinitz, Gross [18] also found a shorter proof for Theorem 7 that is reminiscent of Steinitz’s method. His approach is again based on the rediscovered Steinitz lemma, yet it yields only the weaker constant $S(B_{2}^{d})\leq 2^{d}-1$, which is an inevitable consequence of the induction dimension technique he applies. He also provides a geometric reformulation of Theorem 1: given any closed polygonal path in $\mathbb{R}^{d}$ starting at the origin with side lengths not exceeding 1, it is possible to rearrange the order of the side segments so that the resulting polygonal path does not leave the ball of radius $C$. This later led to the alternate title “polygonal confinement theorem” for Theorem 1 (cf. Rosenthal [29]).
Gross was not the last one to rediscover the Steinitz lemma. In 1931, Bergström published two papers on the topic. In the first [8], he gives an alternative proof for Theorem 7. The crux of his proof is again Theorem 1, which he considers to be of interest on its own. His proof again utilizes induction on the dimension and provides the estimate $S(B_{2}^{d})=\sqrt{(4^{d}-1)/3}$. His second article [9] concentrates solely on the Steinitz lemma. He proves that for $d=2$, the optimal constant is $S(B_{2}^{2})=\sqrt{5}/2$, and formulates Conjecture 4.
In 1936, Hadwiger [19] became the first one to study the Steinitz lemma (which he attributes to Gross and Bergström) for series in general inner product spaces: he manages to bound the norm of partial sums in terms of the number of vectors $n$. His attention then turned to the extension of the Lévy-Steinitz theorem to abstract Hilbert spaces [20] and finite-dimensional vector spaces [21].
Returning to the Euclidean case, Damsteeg and Halperin [14] showed that $S(B_{2}^{d})\geq\frac{1}{2}\sqrt{d+6}$ which implies that an $O(\sqrt{d})$ bound as conjectured by Bergström would be asymptotically optimal (this is also shown by considering the vertex set of a centered regular simplex). Independently of [9], they also determined the exact value of $S(B_{2}^{2})$.
In 1954, Behrend entered the scene [6] and by a refinement of Steinitz’s original method strengthened the estimate to $S(B_{2}^{d})<d$ for every $d\geq 3$. He also showed that $S(B_{2}^{3})\leq\sqrt{5+2\sqrt{3}}$.
Much of the above information had been blocked by the iron curtain. Although Theorem 1 is noted to be a ‘well-known lemma of Steinitz’ [17], Kadets [24] only rediscovered Bergström’s estimate $S(B_{2}^{d})=\sqrt{(4^{d}-1)/3}$ in 1953. This led Belov and Stolin [7] to give a near-optimal algorithm for the flow shop problem. Twenty years later, by a linear algebraic approach, Sevast’yanov [31] re-proved the estimate $S(B_{2}^{d})\leq d$ and showed that this bound holds for all symmetric norms on $\mathbb{R}^{d}$. This led to improved guarantees for the flow shop problem. His proof was further simplified and extended: in 1980, Grinberg and Sevast’yanov [17] gave an elegant, streamlined argument using linear dependencies, which yields the strongest general result, Theorem 3, for not necessarily symmetric norms.
Meanwhile in Hungary, independently of the work in the USSR, Fiala [15] rediscovered the connection between the Steinitz problem and the flow shop problem and re-proved the latter for the case $d=2$. Inspired by his work, Bárány [4] proved the slightly weaker bound $S(B_{2}^{d})\leq 3d/2$, which in turn gives a much more efficient algorithm for the flow shop problem.
Besides estimates for the Euclidean case, many works study the question for other norms. The most important case is $\ell_{\infty}$: Similarly to Conjecture 4, $S(B_{\infty}^{d})=O(\sqrt{d})$ is expected. A lower bound of $\Omega(\sqrt{d})$ was shown using a construction of Hadamard matrices [10], but, similarly to the Euclidean case, no $o(d)$ bound is known. The case $d=2$ was studied by Bárány and Grinberg [11], who proved that $S(B_{\infty}^{2})\leq C$ for a universal constant $C\geq 0$, with the further condition that the first coordinates of each partial sum lie in $[-1,1]$. Banaszyczyk completely classified this case by proving that $S(B_{\infty}^{2})=3/2$ [3], and he also showed [1] that the Steinitz constant of 2-dimensional normed planes does not exceed $3/2$. Furthermore, he gave [2] an alternative proof for the estimate $S(B_{2}^{2})\leq\sqrt{5}/2$ that was established by Bergström half a century ago. Banaszczyk and Sevast’yanov showed a constructive bound of $O(d^{2})$ and a non-constructive bound of $O(d^{3/2})$ for $S(B_{\infty}^{d})$ using techniques specific to the maximum norm [30].
Still in the 1980s, Halperin [22] applied a variation of Lévy’s method to obtain an elementary proof of Theorem 7. In the course of the proof he uses the Grinberg-Sevastyanov variant of the Steinitz lemma, called here as the ‘Polygon Rearrangement Theorem’. He also studies the question in $L_{p}$ and $\ell_{p}$ spaces for $0<p\leq\infty$. In an expository article, Rosenthal [29] presented the Lévy-Steinitz theorem along the Gross-Steinitz approach. He cites the Steinitz lemma as the ‘Polygonal Confinement Theorem’ with the weak bound $S(B_{2}^{d})\leq O(2^{d})$, apparently unaware of its stronger forms.
A possible approach for attacking Conjecture 4 is via Chobanyan’s transference theorem, which gives an explicit connection between the Steinitz constant and the sign-sequence constant $E(B)$, where one is asked to assign signs to vectors of a (potentially infinite) sequence of vectors selected from $B$ so that all partial sums are bounded by $E(B)$. Chobanyan’s result [13, 12] shows that $S(B)\leq E(B)$ [12]. For further information about the sign-sequence constant, see the survey article of Bárány [10].
This brief summary is too short for listing further related topics in detail, such as various extensions of the Lévy-Steinitz theorem to infinite-dimensional spaces (see e.g. [25]), or colorful versions of the Steinitz lemma (see [27] and [5]). Questions on vector sums have a vast literature: the bibliography of Halperin and Ando [23] lists 257 works up to 1988, with at least a hundred related articles having been published since then.
3. Notational Conventions
Throughout the paper, $\log x$ refers to the natural logarithm. For any $10<t<$ 1 and unit vector $u\in S^{d-1}$, we define (see Figure 1):
  • The spherical cap of height $(1-t)$ centered at $u$,
    $C_{t}(u):=\big{\{}v\in S^{d-1}:\ \langle v,u\rangle\geq t\big{\}}.$
  • The ball-cone of height $t$ centered at $u$,
    $K_{t}(u):=\Big{\{}v\in B_{2}^{d}:\ \Big{\langle}\frac{v}{|v|},u\Big{\rangle}\geq t\Big{\}}=\mathrm{conv}\,(C_{t}(u)\cup\{0\});$
Figure 1. Depiction of $C_{t}(u)$ and $K_{t}(u)$.
Let $\kappa_{d}:=\mathrm{Vol}_{d}(B_{2}^{d})=\pi^{d/2}/\Gamma(\nicefrac{{d}}{{2}}+1)$ denote the volume of the Euclidean unit ball and let $\sigma(.)$ be the normalized surface area measure on $S^{d-1}$. We will need to estimate the measure of spherical caps: to that end, we define
$\sigma_{t}:=\sigma(C_{t}(u))$
for an arbitrary $u\in S^{d-1}$ – note that $\sigma_{t}$ is independent of $u$.
4. No short vectors
In this section, we prove the main result of the article.
Proof of Theorem 6.
Take $10\leq\varepsilon<$ 1 as in the statement of Theorem 6, and fix $10<t<$ 1, whose value we will specify later. Suppose that we are given a finite vector set $V\subset B_{2}^{d}$ with $\Sigma(V)=0$; our goal is to order $V$ in such a way that all partial sums are bounded by the right-hand side of (4). The first step of the proof is to partition $V$ as
$V=\Big{(}\dot{\bigcup}_{\alpha\in A}V_{\alpha}\Big{)}\ \dot{\bigcup}\ R,$
where $A$ is an index set of cardinality $m$, satisfying the following properties:
  • (i)
    For each $\alpha\in A$, there exists $u_{\alpha}\in S^{d-1}$ such that $V_{\alpha}\subset K_{t}(u_{\alpha})$
  • (ii)
    For each $\alpha\in A$,
    $\frac{1}{\varepsilon}-1\leq|\Sigma(V_{\alpha})|\leq\frac{1}{\varepsilon}$
  • (iii)
    For any $u\in S^{d-1}$, and any subset $S\subseteq R$,
    $\big{|}\Sigma(S\cap K_{t}(u))\big{|}<1/\varepsilon.$
We define the sets $V_{\alpha}$ via the following process: initialize $R:=V$. As long as there exists $u\in S^{d-1}$ and $S\subseteq R$ such that
(5) $\Big{|}\Sigma(S\cap K_{t}(u))\Big{|}\geq\frac{1}{\varepsilon}-1,$
then we select any subset $V_{\alpha}\subset S\cap K_{t}(u)$ so that
(6) $\frac{1}{\varepsilon}-1\leq\big{|}\Sigma(V_{\alpha})\big{|}\leq\frac{1}{\varepsilon}.$
Such a set is given e.g. by taking any minimal (with respect to containment) subset $T\subset(S\cap K_{t}(u))$ that satisfies $|\Sigma(T)|\geq\frac{1}{\varepsilon}-1$. The family over which we minimize is nonempty by (5), whereas the triangle inequality guarantees that the upper bound in (6) holds as well.
We then update $R:=V\setminus V_{\alpha}$ and proceed until there is no choice of a vector $u\in S^{d-1}$ and $S\subseteq R$ satisfying (5). Property (iii) then follows as for any $u\in S^{d-1}$ and $S\subseteq R$, we have
$\Big{|}\Sigma(S\cap K_{t}(u))\Big{|}<\frac{1}{\varepsilon}-1<\frac{1}{\varepsilon},$
and properties (i) and (ii) follow immediately by construction.
For every $\alpha\in A$, let $w_{\alpha}:=\varepsilon\cdot\Sigma(V_{\alpha})$, and define $W:=\{w_{\alpha}\}_{\alpha\in A}$. Property (ii) yields that for every $\alpha\in A$,
$1-\varepsilon\leq|w_{\alpha}|\leq 1$
Thus, by the definition of the $\varepsilon$-Steinitz constant, there exists an ordering $w_{1},...,w_{m}$ of $W$ such that for any $j\in[m]$,
(7) $\Big{|}\sum_{i\in[j]}w_{i}-\frac{j}{m}\Sigma(W)\Big{|}\leq S_{\varepsilon}(B_{2}^{d}).$
Further, as $\Sigma(V)=0$ and $V=(\dot{\bigcup}_{\alpha\in A}V_{\alpha})\dot{\bigcup}R$,
(8) $\Sigma(W)=\varepsilon\Sigma_{\alpha\in A}\Sigma(V_{\alpha})=-\varepsilon\Sigma(R),$
hence combining (7) and (8) yields the bound
(9) $\Big{|}\sum_{i\in[j]}w_{i}+\varepsilon\frac{j}{m}\Sigma(R)\Big{|}\leq S_{\varepsilon}(B_{2}^{d}).$
We now order our original collection of vectors $V$ as follows: fix the ordering of $W$ as in (7), and note that the ordering of $\{w_{\alpha}\}_{\alpha\in A}$ induces a matching ordering $V_{1},...,V_{m}$ of the sets $\{V_{\alpha}\}_{\alpha\in A}$. Within each set $V_{i}$ for $i\in[m]$ we order the vectors $v\in V_{i}$ arbitrarily. Finally, in the end, we also order the remaining vectors in $R$ arbitrarily. Along this ordering, two types of partial sums occur:
  • (a)
    $\sum_{i\in[j]}\Sigma(V_{i})+\Sigma(T)$ for $j\leq m-1$ and $T\subset V_{j+1}$ (where $T=\emptyset$ is allowed);
  • (b)
    $\sum_{i\in[m]}\Sigma(V_{i})+\Sigma(S)$ for $S\subset R$.
We need to show that any partial sum of type (a) or (b) satisfies the bound in (4). To this end, we will need the following two lemmas, the proofs of which we defer to Section 5.
Lemma 8.
For any $\alpha\in A$ and any $T\subseteq V_{\alpha}$, and for any $10<t<$ 1,
$\big{|}\Sigma(T)\big{|}\leq\frac{1}{\varepsilon t}.$
Lemma 9.
For any $S\subseteq R$, and for any $10<t<$ 1,
$\big{|}\Sigma(S)\big{|}\leq\frac{1}{\varepsilon\sigma_{t}}.$
We now use Lemmas 8 and 9 to complete the proof of Theorem 6. It will be useful to note that by dividing both sides of (9) by $\varepsilon$ and applying the triangle inequality, one has
(10) $\Big{|}\sum_{i\in[j]}\Sigma(V_{i})\Big{|}\leq\frac{j}{k}|\Sigma(R)|+\frac{1}{\varepsilon}S_{\varepsilon}(B_{2}^{d})\leq\frac{1}{\varepsilon\sigma_{t}}+\frac{1}{\varepsilon}S_{\varepsilon}(B_{2}^{d}),$
where we have used that $j\leq m$ and Lemma 9.
We first consider partial sums of type (a). For any $10\leq j\leq m-$ 1 and $T\subseteq V_{j+1}$, combining Lemma 8 and (10) yields
$\Big{|}\sum_{i\in[j]}\Sigma(V_{i})+\Sigma(T)\Big{|}\leq\Big{|}\sum_{i\in[j]}\Sigma(V_{i})\Big{|}+\Big{|}\Sigma(T)\Big{|}\leq\frac{1}{\varepsilon}\left(\frac{1}{\sigma_{t}}+S_{\varepsilon}(B_{2}^{d})+\frac{1}{t}\right),$
where we have interpreted $\Sigma(T)=0$ for the case $T=\emptyset$. Second, we consider any partial sum of type (b). Fix any $S\subseteq R$, and recall that $\sum_{i\in[m]}\Sigma(V_{i})=-\Sigma(R)$. Thus applying Lemma 9 to $R\setminus S\subseteq R$,
$\Big{|}\sum_{i\in[m]}\Sigma(V_{i})+\Sigma(S)\Big{|}=\Big{|}\Sigma(S)-\Sigma(R)\Big{|}=\Big{|}\Sigma(R\setminus S)\Big{|}\leq\frac{1}{\varepsilon\sigma_{t}}.$
We have now shown that every partial sum along the specified ordering has Euclidean norm at most
(11) $\frac{1}{\varepsilon}\left(S_{\varepsilon}(B_{2}^{d})+\frac{1}{t}+\frac{1}{\sigma_{t}}\right).$
We finish the proof via the following estimate on the measure of spherical caps, that we prove in the subsequent section:
Lemma 10.
For every $d\geq 2$, and for
(12) $t=\sqrt{\frac{\log d}{2d}},$
the estimate
(13) $\sigma_{t}\geq\frac{1}{c}\cdot t$
holds with $c=140$.
Thus, setting $t$ as in (12), we have by (13) that
$\frac{1}{t}+\frac{1}{\sigma_{t}}\leq 200\sqrt{\frac{d}{\log d}}$
for every $d\geq 2$, which combined with (11) establishes the desired bound. ∎
We note that by setting $t=\sqrt{\frac{\log d-2\log\log d}{d}}$, Lemma 10 holds with $c=11$ if $d$ is sufficiently large. Consequently, the constant 200 in Theorem 6 can be improved to 12. As this does not constitute an asymptotic improvement, we decided to simplify the calculations by choosing the value of $t$ specified in (12).
5. Technical lemmas
In this section, we prove Lemmas 8, 9, and 10.
Proof of Lemma 8.
Fix $\alpha\in A$ and $T\subseteq V_{\alpha}$. By property (i) of our construction, there exists $u_{\alpha}\in S^{d-1}$ such that $T\subseteq V_{\alpha}\subset K_{t}(u_{\alpha})$.
Since $K_{t}(u_{\alpha})$ is convex, and $T\subseteq V_{\alpha}\subseteq K_{t}(u_{\alpha})$, we have that
(14) $\Big{\langle}\frac{\Sigma(T)}{|\Sigma(T)|},u_{\alpha}\Big{\rangle}\geq t.$
As $T\subseteq V_{\alpha}$ and $0<\langle v,u_{\alpha}\rangle$ for all $v\in V_{\alpha}$,
$\langle\Sigma(T),u_{\alpha}\rangle\leq\langle\Sigma(V_{\alpha}),u_{\alpha}\rangle\leq\big{|}\Sigma(V_{\alpha})\big{|}\leq\frac{1}{\varepsilon}$
by property (ii) of our construction. Thus, by (14),
$|\Sigma(T)|^{2}\leq\frac{1}{t^{2}}\langle\Sigma(T),u_{\alpha}\rangle^{2}\leq\frac{1}{\varepsilon^{2}t^{2}}.\qed$
Proof of Lemma 9.
Fix an arbitrary subset $S\subseteq R$. The key step of the proof is the observation that for any $u\in S^{d-1}$ and $v\in B_{2}^{d}$, then $v\in K_{t}(u)$ if and only if $u\in C_{t}\big{(}\tfrac{v}{|v|}\big{)}$. To see this, note that
$v\in K_{t}(u)\ \ \iff\ \ \Big{\langle}\tfrac{v}{|v|},u\Big{\rangle}\geq t\ \ \ \iff\ \ \ \ v\in C_{t}\big{(}\tfrac{v}{|v|}\big{)}.$
Thus, for any $v\in B_{2}^{d}$, we have
$v=\frac{1}{\sigma_{t}}\int_{S^{d-1}}v\cdot\chi_{K_{t}(u)}(v)\ \mathrm{d}\sigma(u).$
Therefore,
$|\Sigma(S)|^{2}$ $=\langle\Sigma(S),\Sigma(S)\rangle$
$=\Big{\langle}\Sigma(S),\sum_{v\in S}\frac{1}{\sigma_{t}}\int_{S^{d-1}}v\cdot\chi_{K_{t}(u)}(v)\ \mathrm{d}\sigma(u)\Big{\rangle}$
$=\frac{1}{\sigma_{t}}\int_{S^{d-1}}\Big{\langle}\Sigma(S),\Sigma(S\cap K_{t}(u))\Big{\rangle}\ \mathrm{d}\sigma(u)$
$\leq\frac{1}{\sigma_{t}}\int_{S^{d-1}}|\Sigma(S)|\cdot|\Sigma(S\cap K_{t}(u))|\ \mathrm{d}\sigma(u)$
$\leq\frac{|\Sigma(S)|}{\sigma_{t}}\int_{S^{d-1}}\frac{1}{\varepsilon}\ \mathrm{d}\sigma(u)$
$=\frac{|\Sigma(S)|}{\varepsilon\sigma_{t}}$
where we used Property (iii) in the penultimate line. Thus we conclude that $|\Sigma(S)|\leq\frac{1}{\varepsilon\sigma_{t}}$, as desired. ∎
Proof of Lemma 10.
We will use the following bound of Gautschi [16]: for $x>0$ and $10<\lambda<$ 1,
$x^{1-\lambda}\leq\frac{\Gamma(x+1)}{\Gamma(x+\lambda)}\leq(x+1)^{1-\lambda}.$
Combining the above inequality for $x=d/2$ and $\lambda=1/2$ with the definition of $\kappa_{d}$ yields
(15) $\frac{(d-1)\kappa_{d-1}}{d\kappa_{d}}=\frac{(d-1)\cdot\pi^{(d-1)/2}\cdot\Gamma(\frac{d}{2}+1)}{d\cdot\pi^{d/2}\cdot\Gamma(\frac{d}{2}+\frac{1}{2})}\geq\frac{\sqrt{d}}{2\sqrt{2\pi}}.$
A standard calculation shows that
$\sigma_{t}=\frac{(d-1)\kappa_{d-1}}{d\kappa_{d}}\int_{t}^{1}(1-x^{2})^{\frac{d-3}{2}}\ \mathrm{d}x.$
Therefore, by (15),
(16) $\sigma_{t}\geq\frac{\sqrt{d}}{2\sqrt{2\pi}}\int_{t}^{1}(1-x^{2})^{\frac{d-3}{2}}\ \mathrm{d}x\geq\frac{\sqrt{d}}{2\sqrt{2\pi}}\int_{t}^{1}(1-x^{2})^{d/2}\ \mathrm{d}x=\frac{\sqrt{d}}{2\sqrt{2\pi}}\int_{t}^{1}f(x)\ \mathrm{d}x,$
where $f(x):=(1-x^{2})^{d/2}$.
For $d\leq 9$, inequality (13) can be verified by directly calculating the above integral; in particular, for $2\leq d\leq 9$, $\sigma_{t}\geq 0.05$ and $t/c\leq 0.004$.
Assume that $d\geq 10$. Then (12) implies that $t>\frac{1}{\sqrt{d-1}}$. Consequently, since $f(x)$ is convex on the interval $\Big{[}\frac{1}{\sqrt{d-1}},1\Big{]}$, we derive that $f(x)$ is convex on $[t,1]$.
Figure 2. Lower bound on surface area integral using convexity
Consider the triangle $T$ defined by the $x$-axis, the tangent line at $f(t)$, and the line $x=t$, as shown in Figure 2. By convexity,
(17) $\int_{t}^{1}f(x)\ \mathrm{d}x\ \geq\mathrm{Area}(T)=\frac{1}{2}\left(-\frac{f(t)^{2}}{f^{\prime}(t)}\right)=\frac{(1-t^{2})^{d/2+1}}{2dt}.$
Since for $d\geq 4$, $t^{2}=\frac{\log d}{2d}<\frac{1}{5}$, by (16) we obtain that
$\sigma_{t}\geq\frac{1-t^{2}}{4\sqrt{2\pi}}\cdot\frac{(1-t^{2})^{d/2}}{\sqrt{d}t}\geq\frac{1}{5\sqrt{2\pi}}\cdot\frac{(1-t^{2})^{d/2}}{\sqrt{d}t}\geq\frac{1}{13}\cdot\frac{(1-t^{2})^{d/2}}{\sqrt{d}t}\,.$
Thus, the proof of (13) boils down to verifying
$(1-t^{2})^{d/2}\geq\frac{13}{c}\sqrt{d}t^{2},$
which, after substituting (12), takes the form
$\left(1-\frac{\log d}{2d}\right)^{d/2}\geq\frac{13}{2c}\cdot\frac{\log d}{\sqrt{d}}.$
After taking logarithms, this is equivalent to
(18) $\frac{d}{2}\log\left(1-\frac{\log d}{2d}\right)\geq-\log\left(\frac{2c}{13}\right)+\log\log d-\frac{1}{2}\log d.$
By using the standard inequality
$\log(1-\varepsilon)>-\frac{\varepsilon}{1-\varepsilon}$
that holds for every $10<\varepsilon<$ 1, and the estimate $1-\frac{\log d}{2d}\geq\frac{5}{6}$ that is valid for $d\geq 10$, we derive that (18) follows from
(19) $\log d\leq\frac{5}{3}\left(\log d-2\log\log d+2\log\left(\frac{2c}{13}\right)\right).$
Let $d_{0}:=4\cdot 10^{7}$. For $d\geq d_{0}$, $\log\log d<\frac{1}{5}\log d$ holds and implies (19). Finally, for $d\leq d_{0}$, $\log\log d<3$ is valid, and since $\log\left(\frac{2c}{13}\right)>3$, (19) holds trivially. ∎
6. Acknowledgements
We are indebted to Imre Bárány for the fruitful discussions and for providing us with reference [23].